Nice Corporation produces and sells a single product. Data concerning that product appear below:

Per Unit Percent of Sales
Selling price $ 290 100 %
Variable expenses 58 20 %
Contribution margin $ 232 80 %

Fixed expenses are $210,000 per month. The company is currently selling 2,000 units per month.

Required:

Management is considering using a new component that would increase the unit variable cost by $64. Since the new component would improve the company's product, the marketing manager predicts that monthly sales would increase by 700 units. What should be the overall effect on the company's monthly net operating income of this change if fixed expenses are unaffected?
Answer #1
Current Proposed
Unit Sales 2000 (2000+700)=2700
Sales (2000*290)=580,000 (2700*290)=783,000
Variable expenses (2000*58)=116,000 2700*(58+64)=329400
Contribution margin 464,000 453600
Fixed expense 210,000 210,000
Net operating income 254,000 243,600

Hence decrease in Net operating income

=254,000-243,600=$10400.